OpenStudy (anonymous):

solve the system of equations by substitution 3x-y=2 6x=2y+4

hero (hero):

Isolate y in the first equation 3x - y = 2: 3x - 2 = y Substitue y = 3x - 2 into the second equation 6x = 2y + 4 : 6x = 2(3x - 2) + 4

hero (hero):

Solve for x. Then solve for y. Express the solution in the form (x,y).

OpenStudy (anonymous):

ok. i had 6x=6x-4+4, then 6x/6x= 1 is that ok.

OpenStudy (ranga):

6x = 6x - 4 + 4 6x = 6x Whenever you get identical terms on both sides of the equation you can conclude there are INFINITE number of solutions to this system of equations. The two equations represent one and the same line and therefore they share infinite number of common points and therefore they have infinite number of solutions.